# Multi-height analysis of rational points of toric varieties

### v1

@@ 3.4. Conclusion
Let $\epsilon\in]0,1-\frac{1}{f}[$. Then we have:
$\sharp(X(\mathbf{Q}))_{h\in\mathrm{D}_{B}}\underset{B\rightarrow+\infty}{=}\nu(\mathrm{D}_{1}).\tau(X).B^{\langle\omega_{X}^{-1},u\rangle}\left(1+O\left(B^{-(1-\frac{1}{f}-\varepsilon).\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}\right)\right)$
where $\nu$ is the measure over $\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$ defined as in 1.8.
@@
$\left|(\sharp X(\mathbf{Q})_{h\in\mathrm{D}_{B}}-\left(\prod\limits_{p}\omega_{\mathcal{T},p}(\mathcal{T}(\mathbf{Z}_{p}))\right).\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}_{1})B^{\langle\omega_{X}^{-1},u\rangle}\right|\leqslant C.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{(1-\frac{1}{f}-\varepsilon)\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}.$

### v2

@@ 4.4. Conclusion
Let $\epsilon\in]0,1-\frac{1}{l}[$. Then we have:
$\sharp(X(\mathbf{Q}))_{h\in\mathrm{D}_{B}}\underset{B\rightarrow+\infty}{=}\nu(\mathrm{D}_{1}).\tau(X).B^{\langle\omega_{X}^{-1},u\rangle}\left(1+O\left(B^{-(1-\frac{1}{l}-\varepsilon).\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}\right)\right)$
where $\nu$ is the measure over $\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$ defined as in 2.9.
@@
$\left|(\sharp X(\mathbf{Q})_{h\in\mathrm{D}_{B}}-\left(\prod\limits_{p}\omega_{\mathcal{T},p}(\mathcal{T}(\mathbf{Z}_{p}))\right).\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}_{1})B^{\langle\omega_{X}^{-1},u\rangle}\right|\leqslant C.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{(1-\frac{1}{l}-\varepsilon)\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}.$